sir complex number in this question suppose that 3*-8/9+*=3*+8/3+3 in solution in the value of *=59/9 i known but second value 6 5/9 this value i dont know pls help me
Answer by Alan3354(31518)

3*-8/9+*=3*+8/3+3
-------------
* is used for multiplication.
Use x, or n, for something else for the variable.
3x-8/9+x = (3x+8)/3 + 3
-----------
This could be
(3x-8)/9 + x = (3x+8)/3 + 3 or
(3x-8)/(9+x) = (3x+8)/3 + 3 or something else.
Use parentheses to clarify it.